When a ball naturally floats in a pool filled with mercury (density 13600 kg/m° the buoyancy force on the ball is 2224 N. If a force of 1200 N is required to force the ball

completely under water, what is the volume of the ball?


1
Expert's answer
2022-06-03T15:25:34-0400
"g\rho V=mg+F"

"V=\frac{mg+F}{g\rho}=\frac{2224+1200}{9.8*13600}=2.57*10^{-2}\rm m^3"